A man takes 1.5 minutes standing on a moving escalator to reach the top. If escalator is stationary he takes 2.5 minutes to walk up. Then time taken to reach the top when he walk up on moving escalator is nearly
 

Grade:

1 Answers

Khimraj
5 years ago
Let the distance from the ground to the top of escalator = s
speed of the person by walk = s/t2
speed of the escalator = s/t1
effective speed of the person if he/she walks on a moving escalator 
    = s/t1 + s/t2   = s(t1+t2) / (t1 * t2)
Time to reach the top:  s  / speed =  t1 * t2 / (t1 + t2)
here t1 = 1.5
t2 = 2.5
so time taken = 1.5*2.5/4 = 0.9375 minutes
Hope it clears......................